Company A offers a starting salary of $28,000 with a raise of $3,000 each year. Company B offers a starting salary of $36,000 with a raise of $2,000 per year. Company C offers a starting salary of $18000 a year. After how many years would the salaries for companies A and B be the same? How much of a raise would company C have to offer to equal the salaries of Companies A and B in the year in which the salaries of those two companies are the same?
Answer by addingup(3677)

28,000+3,000x= 36,000+2,000x
1,000x= 8,000
x= 8 years
----------------------------
52,000-18,000= 34,000; 34,000/8= annual raise company C would have to offer